OpenStudy (rockets31):

how to simplify the square root of 20

OpenStudy (nenadmatematika):

=2 square roots of 5

OpenStudy (anonymous):

\[\sqrt{20}=\sqrt{2*2*5}=2\sqrt{5}\]

OpenStudy (anonymous):

Okay so steps to simply would be: First, we have a number: \[\sqrt{20}\] So, check how can you factor 20 20=2*2*5 right? So, \[\sqrt{20}=\sqrt{2*2*5}\] Since 2 is repeating (2x2x5-2 is 2 times) We take it out, since \[\sqrt{2*2*5}=\sqrt{2*2}*\sqrt{5}\] Which is same as: \[\sqrt{2*2*5}=\sqrt{4}*\sqrt{5}\] \[\sqrt{4}*\sqrt{5}=2\sqrt{5}\] Which is the final answer.